OpenStudy (steve816):

Please help, I'm stuck at 4 AM doing math!

OpenStudy (steve816):

Solve in the complex number system: \[x^2+3=x\]

OpenStudy (zzr0ck3r):

\(x^2-x+3=0\implies (x-\dfrac{1}{2})^2=-3+(\dfrac{1}{2})^2\\(x-\dfrac{1}{2})^2=\dfrac{-11}{4}\)

OpenStudy (zzr0ck3r):

Can you complete here?

OpenStudy (steve816):

Is completing the square the only method or was I allowed to use the quadratic formula?

OpenStudy (steve816):

I can finish off from here, thanks for the help :)

OpenStudy (zzr0ck3r):

sure, you can use that. This was just faster...

OpenStudy (zzr0ck3r):

The quadratic equation is derived using the "completing the square" method :)

OpenStudy (steve816):

Oh really? Never really understood how the quadratic equation worked even though i'm in pre-calculus.

OpenStudy (zzr0ck3r):

\[ax^2+bx+c=0\\a(x^2+\frac{b}{a}x)=-c\\a(x+\dfrac{b}{2a})^2=-c+a\dfrac{b^2}{4a^2}\\(x+\dfrac{b}{2a})^2=\frac{-c}{a}+\frac{b^2}{4a^2}\\(x+\dfrac{b}{2a})^2=\dfrac{b^2-4ac}{4a^2}\\(x+\dfrac{b}{2a})=\pm \sqrt{\dfrac{b^2-4ac}{4a^2}}\\(x+\dfrac{b}{2a})=\pm \dfrac{\sqrt{b^2-4ac}}{2a}\\x=\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (steve816):

Holy guacamole! :o *jaw drops*

OpenStudy (wolf1728):

Here's another explanation of deriving the quadratic formula: http://www.1728.org/quadr2.htm

OpenStudy (zzr0ck3r):

pretty much the same way.

OpenStudy (wolf1728):

steve - did you check that answer?

OpenStudy (wolf1728):

the (-11/4) answer?

OpenStudy (steve816):

Yup, I got the answer right.

OpenStudy (anonymous):

The answer is (-11/4)

OpenStudy (steve816):

that wasn't the answer... @lauren.ascough

OpenStudy (wolf1728):

I don't get that answer a=1 b=-1 c=3 right?

OpenStudy (anonymous):

Yeah

OpenStudy (wolf1728):

x = 1 +- sq root(1 -4*1*3) / 2 x = [1 +- sq root(-11)] / 2

OpenStudy (steve816):

yeah, that is what I got.

OpenStudy (wolf1728):

Okay, it's just that I didn't see that answer printed anywhere.

OpenStudy (steve816):

Thanks for the answer though.

OpenStudy (anonymous):

Im stuck at 6AM playing and modding skyrim Why? BECAUSE IM IN VIRTUAL SCHOOL